OpenStudy (anonymous):

Translate to an equation. Thirteen less than six times a number is equal to 11 more than four times the number. What is the number?

OpenStudy (anonymous):

6x-13=11*4x

OpenStudy (phi):

it would be good to be able to do this 13 less than means -13+

OpenStudy (phi):

six times a number means 6x if we use "x" to stand for "a number"

OpenStudy (phi):

Thirteen less than six times a number becomes -13 + 6x or, switching the order (when you add, you can swap the order) 6x-13

OpenStudy (phi):

is equal to is written as =

OpenStudy (anonymous):

so how is it supposed to look when i put it all together because i got13-6*a=11+4*a

OpenStudy (anonymous):

is that correct?

OpenStudy (phi):

so far, we have Thirteen less than six times a number is equal to 6x-13=

OpenStudy (phi):

it's ok to use "a" instead of "x" but think about this question: 1 less than 3? the answer is (I hope it is obvious) is 2. to get 2 we do 3 - 1 or -1 + 3 I translate 3 less than x as -3 + x (or x-3) and 3 more than x as 3+x (or x+3)

OpenStudy (anonymous):

ok

OpenStudy (phi):

11 more than four times the number 11 + 4x so all together we have 6x-13= 4x=11

OpenStudy (anonymous):

so their is two = signs?

OpenStudy (phi):

no, that was a typo 6x-13= 4x+11

OpenStudy (anonymous):

i typed it in on my hw online and it says its wrong!

OpenStudy (phi):

The words translate into that equation. You should practice that translation. The next step is to use algebra to solve for x. that is a different skill

OpenStudy (phi):

The question is What is the number? we are not done yet. can you solve for x? in 6x-13= 4x+11 I would add -4x to both sides and then add 13 to both sides

OpenStudy (anonymous):

6x=11 right?

OpenStudy (phi):

let's start with the equation (which you get from translating the words to an equation) 6x-13= 4x+11 add -4x to both sides 6x - 4x - 13 = 4x - 4x + 11 on the right side of the = sign, you have 4x - 4x. that is 0 6x -4x - 13 = 0 + 11 on the right side 0+11 is 11 6x -4x - 13 = 11 on the left side you have 6 x's take away 4 x's. How many x's do you have ?

OpenStudy (anonymous):

1

OpenStudy (phi):

if you don't like x's, how about spoons? 6 spoons take away 4 spoons. How many spoons do you have ?

OpenStudy (anonymous):

2

OpenStudy (phi):

it works the same for x's. 6x's take away 4 x's is 2 x's we write that 6x - 4x = 2x that means 6x -4x - 13 = 11 simplifies to 2x -13 = 11

OpenStudy (phi):

now add +13 to both sides of 2x -13 = 11

OpenStudy (anonymous):

2x=24

OpenStudy (anonymous):

x=12 right?

OpenStudy (phi):

exactly

OpenStudy (phi):

If we use 12 as the number in Thirteen less than six times 12 is equal to 11 more than four times 12 it should work out.